OpenStudy (anonymous):

Monica spent 6/7 hours listening to tapes of Beethoven and Brahms. She spent 1/3 hours listening to Beethoven. How many hours were spent listening to Brahms.

OpenStudy (anonymous):

\[timeBrahms = 6/7 - 1/3\] \[timeBrahms = 18/21 - 7/21\] \[timeBrahms = 11/21\]